OpenStudy (anonymous):

260% write as a fraction - reduce if necessary

OpenStudy (anonymous):

\[ \frac{13}{5} \]

OpenStudy (chriss):

260/100 = 26/10 = 13/5

OpenStudy (anonymous):

2.6/1=26/10=13/5